OpenStudy (anonymous):

y.δ²y/δxδy+2.δu/δx=0 w.r.t u(x,1)=x²+1, u(0,y)=e^y

OpenStudy (anonymous):

I had my own work to do, lol. I suppose the question's really asking to determine \(u(x,y)\). I'm thinking it through.

OpenStudy (anonymous):

lol @Ishaan94 now it looks like I'm talking to myself :P

OpenStudy (anonymous):

Sorry :/

OpenStudy (anonymous):

I think I'll live ;D

OpenStudy (anonymous):

Darn, this is either difficult, or I'm not reading the partials right. Whichever one it is, @TuringTest @FoolForMath @satellite73 @myininaya . See? I'm a part of the problem solving process. :P I get others to do it instead.

OpenStudy (anonymous):

@ayaz110 Could you clarify your question? http://www.codecogs.com/latex/eqneditor.php

OpenStudy (anonymous):

partial derivative with respect to --------

OpenStudy (amistre64):

u(1,0) should get us 1,e

OpenStudy (amistre64):

Ux(x,1) = f'(x,y) = x^2 + y U(x,y) = 1/3 x^3+ yx + g(x,y) Uy = g'(x,y) = x+e^y Uy(0,y) = e^y

OpenStudy (amistre64):

something like this i think, notations bad tho but it helped me focus :)

OpenStudy (amistre64):

U(x,y) = 1/3 x^3 + xy + e^y

OpenStudy (amistre64):

only thing to do is take the required derivatives and see if it fits

OpenStudy (amistre64):

y.δ²y/δxδy+2.δu/δx=0 is that dy/dxdy a typo?

OpenStudy (amistre64):

2.Ux = 2x^2 +2y y Uyx = yx those dont match up to zero tho

OpenStudy (amistre64):

oh well, i think itll be someting similar to that

OpenStudy (anonymous):

I ran into a similar problem, I think there's a typo?

OpenStudy (anonymous):

I'm glad someone else ran into this problem; it confirms that I'm not just derping.